OpenStudy (xxstudyingstuffxx):

Find the percent of the number. 5% of 800 A. 4000 B. 400 C. 40 D. 0.40

OpenStudy (aajc10):

D

OpenStudy (ijlal):

Its C :)

OpenStudy (xxstudyingstuffxx):

What are the steps?

OpenStudy (lukeskylicker):

To find the percentage of a number is quite simple.

OpenStudy (aajc10):

i bet it is d

OpenStudy (lukeskylicker):

if you want to find 5% of 800 all you have to do is multiply 800 by .05

OpenStudy (aajc10):

i guarantee THAT IT IS D

OpenStudy (aajc10):

my bad it is C

OpenStudy (xxstudyingstuffxx):

800 x 0.5 = 400

OpenStudy (ijlal):

5% means 5/100 so 5% of 800 is 800 times 5 and divide by 100

OpenStudy (lukeskylicker):

no

OpenStudy (lukeskylicker):

.05 not .5 that would give you 50%

OpenStudy (xxstudyingstuffxx):

Oh. Do you always put the percentage in the hundredths place, or just one-digit percentages?

OpenStudy (lukeskylicker):

If it's single digit put a 0 in front. Otherwise just type it out as a standard decimal. (or to put it annother way if you have 23% do .23 but if you have 4% do .04)

OpenStudy (gottennis121):

Here's how you do it. What is 5% of 100? There is 800, so multiply 0.5 x 8, and there's your answer.

OpenStudy (xxstudyingstuffxx):

@ijlal is also right. 800 x 5 / 100 is also 40.

OpenStudy (gottennis121):

Yes, so 40 is your answer. :)

OpenStudy (xxstudyingstuffxx):

800 x 5 / 100 = 800 / 0.05 = 40.

OpenStudy (xxstudyingstuffxx):

?

OpenStudy (gottennis121):

To me that's the complicated way, but yeah, that works

OpenStudy (lukeskylicker):

800 x .05 If you divide by a decimal the number gets bigger.

OpenStudy (ijlal):

800 x 5/100 = 800 x 0.05 not 800/0.05

OpenStudy (xxstudyingstuffxx):

800 x 0.5 is 400, @ijlal.

OpenStudy (xxstudyingstuffxx):

800 x 0.05 is 40.

OpenStudy (gottennis121):

Guys, you're making it so complicated... 5% of 100 = ? Use that answer, mulitply time 8, since it's 800. There's your answer

OpenStudy (xxstudyingstuffxx):

Ok, thankyou. Can everybody give eachother a medal. I can't give more than one.

OpenStudy (lukeskylicker):

@gottennis121 What if you have significantly larger numbers that don't go into 100 evenly or numbers smaller than 100 for that matter. (this is genuine question)

OpenStudy (ijlal):

@XxStudyingStuffxX bro 0.05 and 0.5 are different things 50/100 is 0.5 but 5/100 is 0.05 so 5% of 800 is 800 times 5 divide by 100 simple :)

OpenStudy (xxstudyingstuffxx):

Thats why I said you were right.

OpenStudy (xxstudyingstuffxx):

They are just two different ways of doing it.

OpenStudy (gottennis121):

I use that method for that type of stuff that goes into 100. People criticize the way I work things out, but I still get the say answer. :)

OpenStudy (lukeskylicker):

That's still a pretty limited way of getting an answer. That's like using one method that only works if there is a 2 in the equation.

OpenStudy (gottennis121):

I have other ways for different sitiuations.